# DF Louvain: Fast Incrementally Expanding Approach for Community Detection on Dynamic Graphs

### v2

@@ 3.4.2. Delta-screening (DS) approach
Delta-screening, proposed by Zarayeneh et al. (Zarayeneh and Kalyanaraman, 2021), is a dynamic community detection approach that uses modularity-based scoring to determine an approximate region of the graph in which vertices are likely to change their community membership. Figure 1(b) presents a high-level overview of the vertices (and communities), linked to a single source vertex $i$, that are identified as affected using the DS approach in response to a batch update involving both edge deletions and insertions. In the DS approach, Zarayeneh et al. first separately sort the batch update consisting of edge deletions $(i,j)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ by their source vertex-id. For edge deletions within the same community, they mark $i$’s neighbors and $j$’s community as affected. For edge insertions across communities, they pick the highest modularity changing vertex $j*$ among all the insertions linked to vertex $i$ and mark $i$’s neighbors and $j*$’s community as affected. Edge deletions between different communities and edge insertions between the same community are unlikely to affect the community membership of either of the vertices or any other vertices and hence ignored. The affected vertices identified by the DS approach apply to the first pass of Louvain, and the community membership of each vertex is initialized at the start of the community detection algorithm to that obtained in the previous snapshot of the graph. We note that the DS approach, is not guaranteed to explore all vertices that have the potential to change their membership (Zarayeneh and Kalyanaraman, 2021).
@@ 4.1. Our Dynamic Frontier (DF) approach
When a vertex $i$ changes its community during the community detection algorithm (shown by moving $i$ from its original community in the center to its new community on the right), we mark all its neighbor vertices $j\in J_{i}$ as affected, as shown in Figure 1(c) (highlighted in yellow), and mark $i$ as not affected. To minimize unnecessary computation, we also mark an affected vertex $i$ as not affected even if $i$ does not change its community. We call this as the vertex pruning optimization (Ozaki et al., 2016; Ryu and Kim, 2016; Shi et al., 2021; Zhang et al., 2021). The process is akin to a graph traversal and continues until the community assignments of the vertices have converged.
@@
We apply the DF approach to the first pass of Louvain algorithm, as with the DS approach. In subsequent passes all super-vertices are marked as affected and processed according to Louvain. This takes less than $14\%$ of total time, so we don’t use the DF approach to find affected super-vertices. The psuedocode for our parallel DF Louvain is given in Algorithm 1, with its explanation in Section 4.3. Similar to our parallel ND/DS Louvain, it utilizes auxiliary information.
@@ 4.3. Our DF Louvain implementation
In the algorithm, we first identify an initial set of affected vertices, whose communities may directly change due to the batch updates, by marking them in the flag vector $\delta V$. We do this by marking the endpoints of edge deletions $\Delta^{t-}$ which lie in the same community (lines 14-15), and by marking the endpoints of edge insertions $\Delta^{t+}$ which lie is disjoint communities (lines 16-17). We then define three lambda functions for the Louvain algorithm, isAffected() (lines 18-19), inAffectedRange() (lines 20-21), and onChange() (lines 22-23), which indicate that a set of vertices are (initially) marked as affected, that all vertices in the graph can be incrementally marked as affected, and that the neighbors of a vertex are marked as affected if it changes its community membership, respectively. Note that the set of affected vertices will expand automatically due to vertex pruning optimization used in our Parallel Louvain algorithm (Algorithm 4). Thus, onChange() reflects what the DF approach would do in the absence of vertex pruning. Further, unlike existing approaches, we leverage $K^{t-1}$ and $\Sigma^{t-1}$, alongside the batch updates $\Delta^{t-}$ and $\Delta^{t+}$, to efficiently compute $K^{t}$ and $\Sigma^{t}$ required for the local-moving phase of the Louvain algorithm (line 26). These lambda functions and the total vertex/edge weights are then utilized to execute the Louvain algorithm and obtain the updated community assignments $C^{t}$ (line 28). Finally, we return $C^{t}$, alongside $K^{t}$ and $\Sigma^{t}$ (line 29).
@@ 5.1.1. System
For our experiments, we use a server that has an x86-based 64-bit AMD EPYC-7742 processor. This processor has a clock frequency of $2.25$ GHz and $512$ GB of DDR4 system memory. Each core has an L1 cache of $4$ MB, an L2 cache of $32$ MB, and a shared L3 cache of $256$ MB. The machine uses Ubuntu 20.04.
@@ 5.1.2. Configuration
We use 32-bit unsigned integer for vertex IDs, 32-bit floating point for edge weights, but use 64-bit floating point for hashtable values, total edge weight, modularity calculation, and all places where performing an aggregation/sum of floating point values. Affected vertices are represented with an 8-bit integer vector. Computing the weighted degree of each vertex, the local moving phase, and aggregating the edges for the super-vertex graph, employ OpenMP’s dynamic schedule with a chunk size of $2048$ for dynamic workload balancing among threads. We set the iteration tolerance $\tau$ to $10^{-2}$, the tolerance drop per pass $TOLERANCE\_DECLINE\_FACTOR$ to $10$ (threshold-scaling optimization), maximum number of iterations per pass $MAX\_ITERATI$ $ONS$ to $20$, and the maximum number of passes $MAX\_PASSES$ to $10$. Further, we set the aggregation tolerance $\tau_{agg}$ to $0.8$ for large (static) graphs with generated random batch updates, but keep it disabled, i.e., set $\tau_{agg}$ to $1$, for real-world dynamic graphs. Unless mentioned otherwise, we execute all parallel implementations with a default of $64$ threads (to match the number of cores available on the system). We use GCC 9.4 and OpenMP 5.0 (OpenMP Architecture Review Board, 2018) for compilation.
@@ 5.1.4. Batch generation
For experiments involving real-world dynamic graphs, we first load $90\%$ of each graph from Table 2, and ensure that all edges are weigthed with a default weight of $1$, and that they are undirected by adding the reverse edges. Subsequently, we load $B$ edges in $100$ batch updates. Here, $B$ denotes the desired batch size, specified as a fraction of the total number of temporal edges $|E_{T}|$ in the graph, and ensure that the batch update is undirected. For experiments on large graphs with random batch updates, we take each base graph from Table 3 and generate random batch updates (Zarayeneh and Kalyanaraman, 2021) comprising an $80\%:20\%$ mix of edge insertions and deletions to emulate realistic batch updates, each with an edge weight of $1$. To prepare the set of edges for insertion, we select vertex pairs with equal probability. For edge deletions, we uniformly delete each existing edge. To simplify, we ensure no new vertices are added or removed from the graph. The batch size is measured as a fraction of edges in the original graph, ranging from $10^{-7}$ to $0.1$ (i.e., $10^{-7}|E|$ to $0.1|E|$), with multiple batches generated for each size for averaging. For a billion-edge graph, this amounts to a batch size of $100$ to $100$ million edges. All batch updates are undirected, i.e., for every edge insertion $(i,j,w)$ in the batch update, the edge $(j,i,w)$ is also a part of the batch update. We employ five distinct random batch updates for each batch size, and report average across these runs in our experiments.
@@ 5.1.5. Measurement
We evaluate the runtime of each approach on the entire updated graph, including the local-moving phase, aggregation phase, the initial and incremental marking of affected vertices, convergence detection, and all the necessary intermediary steps, but excluding memory allocation/deallocation time. We assume that the total edge weight of the graphs is known, and can be tracked upon each batch update.
@@
Figure 5. Mean Runtime and Modularity of communities obtained with our multicore implementation of Static, Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) PageRank on real-world dynamic graphs, with batch updates of size $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$. Here, (a) and (b) show the overall runtime and modularity across all temporal graphs, while (c) and (d) show the runtime and modularity for each graph. In (a), the speedup of each approach with respect to Static Louvain is labeled.
@@ 5.2.1. Results on real-world dynamic graphs
Figure 5(a) shows that DF Louvain is, on average, 268×268$\times$268 ×, 195×195$\times$195 ×, and 109×109$\times$109 × faster than Static Louvain on batch updates of size $10^{-5}|E_{T}|$, $10^{-4}|E_{T}|$, and $10^{-3}|E_{T}|$, respectively. In contrast, DS Louvain demonstrates average speedups of 47×47$\times$47 ×, 32×32$\times$32 ×, and 26×26$\times$26 × over Static Louvain for batch update sizes of $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$, while ND Louvain obtains a mean speedup of 25×25$\times$25 ×. DF Louvain is thus, overall, 179×179$\times$179 ×, 7.2×7.2$\times$7.2 ×, and 5.3×5.3$\times$5.3 × faster than Static, ND, and DS Louvain. This speedup is particularly pronounced on the sx-superuser graph, as indicated by Figure 5(c). Regarding modularity, Figures 5(b) and 5(d) illustrate that DF Louvain generally exhibits slightly lower modularity on average compared to ND and DS Louvain but is on par with the modularity obtained by Static Louvain, except for the sx-superuser graph (because it fails to mark certain vertices as affected, likely because they are outlier vertices and were not directly reached from the expanding frontier). This makes the communities obtained with DF Louvain generally acceptable. However, if lower modularity is observed (during intermediate empirical tests), transitioning to our parallel implementation of DS Louvain is advisable.
@@ 6. Conclusion
In conclusion, in this report we presented our Parallel Dynamic Frontier (DF) Louvain algorithm, which given a batch update of edge deletions or insertions, incrementally identifies and processes an approximate set of affected vertices in the graph with minimal overhead. In addition, we use a novel approach of incrementally updating weighted-degrees of vertices and total edge weights of communities, by using them as auxiliary information to DF Louvain. We also present our parallel implementations of Naive-dynamic (ND) (Aynaud and Guillaume, 2010) and Delta-screening (DS) Louvain (Zarayeneh and Kalyanaraman, 2021). On a server with a 64-core AMD EPYC-7742 processor, our experiments show that DF Louvain obtains speedups of 179×179$\times$179 ×, 7.2×7.2$\times$7.2 ×, and 5.3×5.3$\times$5.3 × on real-world dynamic graphs, compared to Static, ND, and DS Louvain, respectively, and is 183×183$\times$183 ×, 13.8×13.8$\times$13.8 ×, and 8.7×8.7$\times$8.7 × faster, respectively, on large graphs with random batch updates. The performance of DF Louvain can be attributed to incremental marking, and marking of fewer vertices. Furthermore, DF Louvain improves performance at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads. DF Louvain is based on one of the fastest implementations of Static Louvain (Sahu, 2023) and thus offers state-of-the-art performance. However, we observe DF Louvain to obtain lower modularity score on the sx-superuser real-world dynamic graph. We thus recommend the reader to use DF Louvain for updating community memberships on dynamic graphs, but switch to DS Louvain if lower modularity is observed.
@@ A.5.4. Overall
Finally, based on Lemmas A.1, A.2, and A.3, we can state the following for DF Louvain.
Theorem A.4.

### v4

@@ 3.4.2. Delta-screening (DS) approach
Delta-screening, proposed by Zarayeneh et al. (Zarayeneh and Kalyanaraman, 2021), is a dynamic community detection approach that uses modularity-based scoring to determine an approximate region of the graph in which vertices are likely to change their community membership. Figure 1(b) presents a high-level overview of the vertices (and communities), linked to a single source vertex $i$, that are identified as affected using the DS approach in response to a batch update involving both edge deletions and insertions. In the DS approach, Zarayeneh et al. first separately sort the batch update consisting of edge deletions $(i,j)\in\Delta^{t-}$ and insertions $(i,j,w)\in\Delta^{t+}$ by their source vertex-id. For edge deletions within the same community, they mark $i$’s neighbors and $j$’s community as affected. For edge insertions across communities, they pick the highest modularity changing vertex $j*$ among all the insertions linked to vertex $i$ and mark $i$’s neighbors and $j*$’s community as affected. Edge deletions between different communities and edge insertions between the same community are unlikely to affect the community membership of either of the vertices or any other vertices and hence ignored. The affected vertices identified by the DS approach apply to the first pass of Louvain algorithm, and the community membership of each vertex is initialized at the start of the community detection algorithm to that obtained in the previous snapshot of the graph. We note that the DS approach, is not guaranteed to explore all vertices that have the potential to change their membership (Zarayeneh and Kalyanaraman, 2021).
@@ 4.1. Our Dynamic Frontier (DF) approach
When a vertex $i$ changes its community during the community detection algorithm (shown by moving $i$ from its original community in the center to its new community on the right), we mark all its neighbor vertices $j\in J_{i}$ as affected, as shown in Figure 1(c) (highlighted in yellow), and mark $i$ as not affected. If $i$ does not change its community, it is also marked as unaffected — this is known as vertex pruning optimization (Ozaki et al., 2016; Ryu and Kim, 2016; Shi et al., 2021; Zhang et al., 2021). The process is akin to a graph traversal and continues until the community assignments of the vertices have converged.
@@
We apply the DF approach to the first pass of Louvain algorithm, as with the DS approach. In subsequent passes all super-vertices are processed, according to the Louvain algorithm. This takes less than $15\%$ of total time, on average, so we don’t use the DF approach to find affected super-vertices. The psuedocode for our parallel DF Louvain is given in Algorithm 1, with its explanation in Section 4.3. Similar to our parallel ND/DS Louvain, it utilizes auxiliary information.
@@ 4.3. Our DF Louvain implementation
In the algorithm, we first identify an initial set of affected vertices, whose communities may directly change due to the batch updates, by marking them in the flag vector $\delta V$. We do this by marking the endpoints of edge deletions $\Delta^{t-}$ which lie in the same community (lines 14-15), and by marking the endpoints of edge insertions $\Delta^{t+}$ which lie in disjoint communities (lines 16-17). We then define three lambda functions for the Louvain algorithm, isAffected() (lines 18-19), inAffectedRange() (lines 20-21), and onChange() (lines 22-23), which indicate that a set of vertices are (initially) marked as affected, that all vertices in the graph can be incrementally marked as affected, and that the neighbors of a vertex are marked as affected if it changes its community membership, respectively. Note that the set of affected vertices will expand automatically due to vertex pruning optimization used in our Parallel Louvain algorithm (Algorithm 4). Thus, onChange() reflects what the DF approach would do in the absence of vertex pruning. Further, unlike existing approaches, we leverage $K^{t-1}$ and $\Sigma^{t-1}$, alongside the batch updates $\Delta^{t-}$ and $\Delta^{t+}$, to efficiently compute $K^{t}$ and $\Sigma^{t}$ required for the local-moving phase of the Louvain algorithm (line 26). These lambda functions and the total vertex/edge weights are then utilized to execute the Louvain algorithm and obtain the updated community assignments $C^{t}$ (line 28). Finally, we return $C^{t}$, alongside $K^{t}$ and $\Sigma^{t}$ (line 29).
@@ 5.1.1. System
For our experiments, we use a server that has an x86-based 64 core AMD EPYC-7742 processor. This processor has a clock frequency of $2.25$ GHz and $512$ GB of DDR4 system memory. Each core has an L1 cache of $4$ MB, an L2 cache of $32$ MB, and a shared L3 cache of $256$ MB. The machine uses Ubuntu 20.04.
@@ 5.1.2. Configuration
We use 32-bit unsigned integer for vertex IDs, 32-bit floating point for edge weights, but use 64-bit floating point for hashtable values, total edge weight, modularity calculation, and all places where performing an aggregation/sum of floating point values. Affected vertices are represented with an 8-bit integer vector. Computing the weighted degree of each vertex, the local moving phase, and aggregating the edges for the super-vertex graph, employ OpenMP’s dynamic schedule with a chunk size of $2048$ for dynamic workload balancing among threads. We set the iteration tolerance $\tau$ to $10^{-2}$, the tolerance drop per pass $TOLERANCE\_DECLINE\_FACTOR$ to $10$ (threshold-scaling optimization), maximum number of iterations per pass $MAX\_ITERATI$ $ONS$ to $20$, and the maximum number of passes $MAX\_PASSES$ to $10$ (Sahu, 2023). Further, we set the aggregation tolerance $\tau_{agg}$ to $0.8$ for large (static) graphs with generated random batch updates, but keep it disabled, i.e., set $\tau_{agg}$ to $1$, for real-world dynamic graphs. Unless mentioned otherwise, we execute all parallel implementations with a default of $64$ threads (to match the number of cores available on the system). We use GCC 9.4 and OpenMP 5.0 (OpenMP Architecture Review Board, 2018) for compilation.
@@ 5.1.4. Batch generation
For experiments involving real-world dynamic graphs, we first load $90\%$ of the temporal edges of each graph from Table 2, and ensure that all edges are weighted with a default weight of $1$, and that they are undirected by adding the reverse edges. Subsequently, we load $B$ edges in $100$ batch updates. Here, $B$ denotes the desired batch size, specified as a fraction of the total number of temporal edges $|E_{T}|$ in the graph, and ensure that the batch update is undirected. For experiments on large graphs with random batch updates, we take each base graph from Table 3 and generate random batch updates (Zarayeneh and Kalyanaraman, 2021) comprising an $80\%:20\%$ mix of edge insertions and deletions to emulate realistic batch updates, each with an edge weight of $1$. To prepare the set of edges for insertion, we select vertex pairs with equal probability. For edge deletions, we remove the desired number of existing edges, with equal probability of selecting each edge. To simplify, we ensure no new vertices are added or removed from the graph. The batch size is measured as a fraction of edges in the original graph, ranging from $10^{-7}$ to $0.1$ (i.e., $10^{-7}|E|$ to $0.1|E|$), with multiple batches generated for each size for averaging. For a billion-edge graph, this amounts to a batch size of $100$ to $100$ million edges. All batch updates are undirected, i.e., for every edge insertion $(i,j,w)$ in the batch update, the edge $(j,i,w)$ is also a part of the batch update. We employ five distinct random batch updates for each batch size, and report average across these runs in our experiments.
@@ 5.1.5. Measurement
We evaluate the runtime of each approach on the entire updated graph, including the local-moving phase, aggregation phase, the initial and incremental marking of affected vertices, convergence detection, and all the necessary intermediary steps, but excluding only the initial memory allocation/deallocation time. We assume that the total edge weight of the graphs is known, and can be tracked upon each batch update.
@@
Figure 5. Mean Runtime and Modularity of communities obtained with our multicore implementation of Static, Naive-dynamic (ND), Delta-screening (DS), and Dynamic Frontier (DF) Louvain on real-world dynamic graphs, with batch updates of size $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$. Here, (a) and (b) show the overall runtime and modularity across all temporal graphs, while (c) and (d) show the runtime and modularity for each graph. In (a), the speedup of each approach with respect to Static Louvain is labeled.
@@ 5.2.1. Results on real-world dynamic graphs
Figure 5(a) shows that DF Louvain is, on average, 268×268$\times$268 ×, 195×195$\times$195 ×, and 109×109$\times$109 × faster than Static Louvain on batch updates of size $10^{-5}|E_{T}|$, $10^{-4}|E_{T}|$, and $10^{-3}|E_{T}|$, respectively. In contrast, DS Louvain demonstrates average speedups of 47×47$\times$47 ×, 32×32$\times$32 ×, and 26×26$\times$26 × over Static Louvain for batch update sizes of $10^{-5}|E_{T}|$ to $10^{-3}|E_{T}|$, while ND Louvain obtains a mean speedup of 25×25$\times$25 ×. DF Louvain is thus, overall, 179×179$\times$179 ×, 7.2×7.2$\times$7.2 ×, and 5.3×5.3$\times$5.3 × faster than Static, ND, and DS Louvain. This speedup is particularly pronounced on the sx-superuser graph, as indicated by Figure 5(c). Regarding modularity, Figures 5(b) and 5(d) illustrate that DF Louvain generally exhibits slightly lower modularity on average compared to ND and DS Louvain but is on par with the modularity obtained by Static Louvain, except for the sx-superuser graph (because it fails to mark certain vertices as affected, likely because they are outlier vertices and were not directly reached from the expanding frontier). This makes the communities obtained with DF Louvain generally acceptable. However, if lower modularity is observed (during intermediate empirical tests), transitioning to our parallel implementation of DS Louvain is advisable. It is also worth noting that ND Louvain produces higher-quality communities than Static Louvain, as it builds upon the community memberships obtained from Static Louvain and further optimizes them (we limit the number of passes for Static Louvain to achieve a large speed gain, even if it means sacrificing some quality).
@@ 6. Conclusion
In conclusion, in this report we presented our Parallel Dynamic Frontier (DF) Louvain algorithm, which given a batch update of edge deletions or insertions, incrementally identifies and processes an approximate set of affected vertices in the graph with minimal overhead. In addition, we use a novel approach of incrementally updating weighted-degrees of vertices and total edge weights of communities, by using them as auxiliary information to DF Louvain. We also present our parallel implementations of Naive-dynamic (ND) (Aynaud and Guillaume, 2010) and Delta-screening (DS) Louvain (Zarayeneh and Kalyanaraman, 2021). On a server with a 64-core AMD EPYC-7742 processor, our experiments show that DF Louvain obtains speedups of 179×179$\times$179 ×, 7.2×7.2$\times$7.2 ×, and 5.3×5.3$\times$5.3 × on real-world dynamic graphs, compared to Static, ND, and DS Louvain, respectively, and is 183×183$\times$183 ×, 13.8×13.8$\times$13.8 ×, and 8.7×8.7$\times$8.7 × faster, respectively, on large graphs with random batch updates. Note that our static implementation of the Louvain algorithm, GVE-Louvain, here referred to Static Louvain, is itself 50×50$\times$50 ×, $\times2$, 20×20$\times$20 ×, and 5.8×5.8$\times$5.8 × faster than Vite, Grappolo, NetworKit Louvain, and cuGraph Louvain (running on NVIDIA A100 GPU) (Sahu, 2023). The performance of DF Louvain can be attributed to incremental marking, and marking of fewer vertices. Furthermore, DF Louvain improves performance at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads. DF Louvain is based on one of the fastest implementations of Static Louvain (Sahu, 2023) and thus offers state-of-the-art performance. However, we observe DF Louvain to obtain lower modularity score on the sx-superuser real-world dynamic graph. We thus recommend the reader to use DF Louvain for updating community memberships on dynamic graphs, but switch to DS Louvain if lower modularity is observed.
@@ A.5.4. Overall
Finally, based on Observations A.1, A.2, and A.3, we can state the following for DF Louvain.
Conclusion A.4.